# Topologizability and Power Boundedness of Convolutions and Toeplitz Operators on Power Series Spaces

### v2

@@ Abstract.
We completely characterize the topologizability and power boundedness of convolution and dual convolution operators on power series spaces. We determine necessary conditions for a Toeplitz operator to be m-topologizable, and power bounded on $\Lambda_{1}(n)$ and $\Lambda_{\infty}(n)$, and consequently on $H(\mathbb{C})$ and $H(\mathbb{D})$.
@@ 1. Introduction
The terminology of topologizable and m-topologizable operators was first formulated by Żelazko in [13], from the perspective of equipping the space of continuous linear operators $\mathcal{L}(E)$, where $E$ is a non-normable locally convex space, with a compatible locally convex topology. In recent years, topologizability, power boundedness, mean ergodicity, and related properties have been extensively explored for continuous linear operators on locally convex spaces, especially Fréchet spaces, see, for instance, [1], [2], [3], [4], [12], [13].
@@ 2. Preliminaries
Żelazko in [15, Proposition 7 and Proposition 12] showed that the sum and the product of two commuting topologizable (respectively, m-topologizable) operators are topologizable (respectively, m-topologizable) too. In Example 13 of [15], Żelazko provides a counterexample showing that the sum of two m-topologizable operators, which do not commute, need not be topologizable.
@@ 3. Toeplitz Operators on Power Series Spaces
We would like to note that, as shown in Theorems 3.1 and 3.2, the entries of the Toeplitz matrix come from the Laurent series coefficients of the symbol $F$ centered at $x=0$.
@@ 3.1. The Convolution Operators $\widehat{T}_{\theta}$ Acting on Power Series Spaces
Since they will be used later, the computations given in Theorem 3.2 and Theorem 3.5 of [6] will be highlighted in the following two propositions.
@@
Proof.
Let $\theta\in\Lambda_{1}(\alpha)$. For every $p,n\in\mathbb{N}$, we have
that is,
$\|\widehat{T}_{\theta}e_{n}\|_{p}\leq\|\theta\|_{2p}\|e_{n}\|_{2p}$ (3.3)
holds for every $n,p\in\mathbb{N}$. Therefore, $\widehat{T}_{\theta}e_{n}\in\Lambda_{1}(\alpha)$ for every $n\in\mathbb{N}$ and for every $p\in\mathbb{N}$
$\sup_{n\in\mathbb{N}}\frac{\|\widehat{T}_{\theta}e_{n}\|_{p}}{\|e_{n}\|_{2p}}=\|\theta\|_{2p}<\infty.$
Hence $\widehat{T}_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ is well defined and continuous by Proposition 2.2. ∎
@@
Proof.
and if $j=2t+1\geq 2n$, we have $t+1\geq n$ and
Therefore, we have
$\alpha_{j}\leq M\alpha_{j-n+1}$
for all $j\geq 2n$. For every $p,n\in\mathbb{N}$, we have
that is,
$\|\widehat{T}_{\theta}e_{n}\|_{p}\leq\|\theta\|_{Mp}\|e_{n}\|_{Mp}$ (3.4)
holds for every $n,p\in\mathbb{N}$. Therefore, $\widehat{T}_{\theta}e_{n}\in\Lambda_{\infty}(\alpha)$ for every $n\in\mathbb{N}$ and for every $p\in\mathbb{N}$
$\sup_{n\in\mathbb{N}}\frac{\|\widehat{T}_{\theta}e_{n}\|_{p}}{\|e_{n}\|_{Mp}}=\|\theta\|_{Mp}<\infty.$
Hence $\widehat{T}_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ is well defined and continuous by Proposition 2.2. ∎
@@ 3.2. The Dual Convolution Operators $\widecheck{T}_{\beta}$ Acting on Power Series Spaces
Since we will make use of them later, the calculations given in Theorem 4.5 and Theorem 4.3 of [6] will be emphasized in the following two propositions.
@@
Proof.
Let $\beta\in(\Lambda_{1}(\alpha))^{\prime}$. Then there exist some $m_{0}\in\mathbb{N}$ and $C_{0}>0$ such that
$|\beta_{n-1}|\leq C_{0}e^{-\frac{1}{m_{0}}\alpha_{n}}$
for all $n\in\mathbb{N}.$ Since $\alpha$ is stable, there exists a $M\in\mathbb{N}$ such that $\alpha_{2t}\leq M\alpha_{t}$ for every $t\in\mathbb{N}$. Assume that $n=2t$ or $n=2t+1$ and $1\leq j\leq n$
$\alpha_{n}\leq M(\alpha_{n-j+1}+\alpha_{j})$ (3.5)
is satisfied for all $n\in\mathbb{N}$ and $j\leq n$. For every $p,n\in\mathbb{N}$, we have
$\begin{split}\|\widecheck{T}_{\theta}e_{n}\|_{p}&=\sum^{n}_{j=1}|\beta_{n-j}|e^{-\frac{1}{p}\alpha_{j}}\leq C_{0}\sum^{n}_{j=1}e^{-\frac{1}{m_{0}}\alpha_{n-j+1}}e^{-\frac{1}{p\alpha_{j}}}\\ &\leq C_{0}\sum^{n}_{j=1}e^{-\frac{1}{\max\{m_{0},p\}}(\alpha_{n-j+1}+\alpha_{j})}\leq C_{0}\sum^{n}_{j=1}e^{-\frac{1}{M\max\{m_{0},p\}}\alpha_{n}}\\ &\leq C_{0}ne^{-\frac{1}{M\max\{m_{0},p\}}\alpha_{n}}\end{split}$
for every $k\in\mathbb{N}$. It follows that there exists a $D_{p}>0$ satisfying
$ne^{-\frac{1}{M\max\{m_{0},p\}}\alpha_{n}}\leq D_{p}e^{-\frac{1}{2M\max\{m_{0},p\}}\alpha_{n}}$ (3.6)
for every $n\in\mathbb{N}$ and this implies that
$\begin{split}\|\widecheck{T}_{\beta}e_{n}\|_{p}&\leq C_{0}ne^{-\frac{1}{M\max\{m_{0},p\}}\alpha_{n}}\\ &\leq C_{0}D_{p}e^{-\frac{1}{2M\max\{m_{0},p\}}\alpha_{n}}=C_{0}D_{p}\|e_{n}\|_{{2M\max\{m_{0},p\}}}\end{split}$ (3.7)
for every $p,n\in\mathbb{N}$. Therefore, $\widecheck{T}_{\beta}e_{n}\in\Lambda_{1}(\alpha_{n})$ for every $n\in\mathbb{N}$ and for every $p\in\mathbb{N}$, there exists $q=2M\max\{m_{0},p\}\in\mathbb{N}$ such that
$\sup_{n\in\mathbb{N}}\frac{\|\widecheck{T}_{\beta}e_{n}\|_{p}}{\|e_{n}\|_{q}}<\infty.$
That is, the operator $\widecheck{T}_{\beta}:\Lambda_{1}(\alpha)\to K(b_{n,k})$ is well defined and continuous by Proposition 2.2. ∎
@@
Proof.
Let $\beta\in(\Lambda_{\infty}(\alpha))^{\prime}$. Then there exist some $m_{0}\in\mathbb{N}$ and $C_{0}>0$ such that
$|\beta_{n-1}|\leq C_{0}e^{m_{0}\alpha_{n}}$
for all $n\in\mathbb{N}.$ For every $p,n\in\mathbb{N}$, the inequalities
$\begin{split}\|\widecheck{T}_{\theta}e_{n}\|_{p}=\sum^{n}_{j=1}|\theta_{n-j}|e^{p\alpha_{j}}\leq C_{0}\sum^{n}_{j=1}e^{m_{0}\alpha_{n-j+1}}e^{p\alpha_{j}}\leq C_{0}e^{m_{0}\alpha_{n}}\sum^{n}_{j=1}e^{p\alpha_{j}}\end{split}$
hold. Since $\Lambda_{\infty}(\alpha)$ is nuclear, $\sum^{\infty}_{n=1}e^{-m_{1}\alpha_{n}}$ is convergent for some $m_{1}\in\mathbb{N}$. It follows that there exists a $D>0$ such that
$\sum^{n}_{j=1}e^{p\alpha_{j}}\leq De^{(p+m_{1})\alpha_{n}}.$ (3.8)
Indeed, this is evident from the following estimates:
Then, we get
for every $n\in\mathbb{N}$. Therefore, $\widecheck{T}_{\beta}e_{n}\in\Lambda_{\infty}(\alpha)$ for every $n\in\mathbb{N}$ and for all $p\in\mathbb{N}$, there exist $q=p+m_{0}+m_{1}\in\mathbb{N}$ such that
$\sup_{n\in\mathbb{N}}\frac{\|\widecheck{T}_{\beta}e_{n}\|_{p}}{\|e_{n}\|_{q}}<\infty,$
that is, $\widecheck{T}_{\beta}$ is well defined and continuous by Proposition 2.2. ∎
@@ 3.3. Products Within Convolution and Dual Convolution Operator Classes
In the proof Proposition of 3.7, we used the nuclearity of $\Lambda_{\infty}(\alpha)$, to show that there exist $m_{1}\in\mathbb{N}$ and $D>0$ such that for all $p,n\in\mathbb{N}$
$\sum^{n}_{j=1}e^{p\alpha_{j}}\leq De^{(p+m_{1})\alpha_{n}}.$
@@
In the proof of Proposition 3.6, we used the nuclearity of $\Lambda_{1}(\alpha)$, to show that for every $k\in\mathbb{N}$, there exists a $D>0$ satisfying
$ne^{-\frac{1}{k}\alpha_{n}}\leq De^{-\frac{1}{2k}\alpha_{n}}$
for all $n\in\mathbb{N}$. Moreover, in the same proof, the stability of $\alpha$ imply the existence of a constant $M\in\mathbb{N}$ such that
$\alpha_{n}\leq M(\alpha_{n-j+1}+\alpha_{j})$
for every $j\leq n$ and $n\in\mathbb{N}$ .
@@
$\widecheck{T}_{\beta}\widecheck{T}_{\theta}=\widecheck{T}_{\psi}\widecheck{T}_{\beta}=\widecheck{T}_{\psi*\beta}.$
@@ 4. Topologizability and Power Boundedness of Convolution Operators on Power Series Spaces
Let $\theta\in\Lambda_{1}(\alpha)$. In the proof of Proposition 3.3, specifically in (3.3), we showed that
@@
$\theta^{*k}:=\underbrace{\theta*\theta*\cdots*\theta}_{k\ \text{times}}$
@@
$\begin{split}\|\theta^{*k}\|_{p}&=\|\widehat{T}_{\theta^{*k}}e_{1}\|_{p}=\|\widehat{T}_{\theta}(\widehat{T}_{\theta^{*(k-1)}}e_{1})\|_{p}=\\ &\leq\|\theta\|_{2p}\|\widehat{T}_{\theta^{*(k-1)}}e_{1}\|_{2p}\leq\dots\leq(\|\theta\|_{2p})^{k}\end{split}$ (4.3)
@@
$\|\widehat{T}^{k}_{\theta}e_{n}\|_{p}\leq(\|\theta\|_{2p})^{k}\|e_{n}\|_{2p}.$ (4.4)
for every $p,k,n\in\mathbb{N}$. This ensures that $\widehat{T}_{\theta}$ is topologizable, even m-topologizable for every $\theta\in\Lambda_{1}(\alpha)$.
@@
The convolution operator $\widehat{T}_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ is topologizable, even m-topologizable for every $\theta\in\Lambda_{1}(\alpha)$.
@@
The proof follows immediately from the inequality (4.4) and Lemma 2.5. ∎
Given that the operators $\widehat{T}_{\theta}$ and $\widehat{T}_{\phi}$ commute and their product satisfies $\widehat{T}_{\theta}\widehat{T}_{\phi}=\widehat{T}_{\theta*\phi}$, we can conclude that their product is always topologizable - and even m-topologizable for every $\theta,\phi\in\Lambda_{1}(\alpha)$.
@@
Let $\theta\in\Lambda_{\infty}(\alpha)$. In the proof of Proposition 3.4, specifically in (3.4), we showed that
@@
holds for every $p,n\in\mathbb{N}$ and here, $M\in\mathbb{N}$ denotes a constant satisfying $\sup_{n\in\mathbb{N}}\frac{\alpha_{2n}}{\alpha_{n}}<M$. In Proposition 3.9, we showed that the product of two convolution operators, say $\widehat{T}_{\theta}$ and $\widehat{T}_{\phi}$, is again a convolution operator $\widehat{T}_{\theta*\phi}$. By using (4.5), we obtain
@@
$\begin{split}\|\theta^{*k}\|_{p}&=\|\widehat{T}_{\theta^{*k}}e_{1}\|_{p}=\|\widehat{T}_{\theta}(\widehat{T}_{\theta^{*(k-1)}}e_{1})\|_{p}\\ &\leq\|\theta\|_{Mp}\|\widehat{T}_{\theta^{*(k-1)}}e_{1}\|_{2p}\leq\dots\leq(\|\theta\|_{Mp})^{k}\end{split}$ (4.7)
@@
$\|\widehat{T}^{k}_{\theta}e_{n}\|_{p}\leq(\|\theta\|_{Mp})^{k}\|e_{n}\|_{Mp}.$ (4.8)
holds for every $p,k,n\in\mathbb{N}$. This ensures that $\widehat{T}_{\theta}$ is topologizable, even m-topologizable for every $\theta\in\Lambda_{\infty}(\alpha)$.
@@
Let $\alpha$ be a stable sequence. The convolution operator $\widehat{T}_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\alpha)$ is topologizable, even m-topologizable for every $\theta\in\Lambda_{\infty}(\alpha)$.
@@
The proof follows immediately from the inequality (4.8) and Lemma 2.5. ∎
We again note that the operators $\widehat{T}_{\theta}$ and $\widehat{T}_{\phi}$ commute and their product satisfies $\widehat{T}_{\theta}\widehat{T}_{\phi}=\widehat{T}_{\theta*\phi}$, then we can conclude that their product is always topologizable - and even m-topologizable for every $\theta,\phi\in\Lambda_{\infty}(\alpha)$.

### v3

@@ Abstract.
We characterize the topologizability and power boundedness of convolution and dual convolution operators on power series spaces. We determine necessary conditions for a Toeplitz operator to be m-topologizable, and power bounded on $\Lambda_{1}(n)$ and $\Lambda_{\infty}(n)$, and consequently on $H(\mathbb{C})$ and $H(\mathbb{D})$.
@@ 1. Introduction
The terminology of topologizable and m-topologizable operators was first formulated by Żelazko in [13], from the perspective of equipping the space of continuous linear operators $\mathcal{L}(E)$, where $E$ is a non-normable locally convex space, with a compatible locally convex topology. In recent years, topologizability, power boundedness, mean ergodicity, and related properties have been extensively explored for continuous linear operators on locally convex spaces, especially Fréchet spaces, see, for instance, [1], [2], [3], [4], [8] [13], [14].
@@ 2. Preliminaries
Żelazko in [16, Proposition 7 and Proposition 12] showed that the sum and the product of two commuting topologizable (respectively, m-topologizable) operators are topologizable (respectively, m-topologizable) too. In Example 13 of [16], Żelazko provides a counterexample showing that the sum of two m-topologizable operators, which do not commute, need not be topologizable.
@@ 3. Toeplitz Operators on Power Series Spaces
We would like to note that, as shown in Theorems 3.1 and 3.2, the entries of the Toeplitz matrix come from the Laurent series coefficients of the symbol $F$ centered at $z=0$.
@@ 3.1. The Convolution Operators $\widehat{T}_{\theta}$ Acting on Power Series Spaces
The following two propositions, 3.3 and 3.4, correspond to special cases of Theorems 3.3 and 3.6 in [6]. For brevity, we omit the proofs and direct the reader to [6] for the full computations.
@@ 3.2. The Dual Convolution Operators $\widecheck{T}_{\beta}$ Acting on Power Series Spaces
The following two propositions, 3.6 and 3.7, correspond to special cases of Theorems 4.6 and 4.4 in [6]. For brevity, we omit the proofs and refer the reader to [6] for the full computations.
@@ 3.3. Products Within Convolution and Dual Convolution Operator Classes
Since $\Lambda_{\infty}(\alpha)$ is nuclear, $\sum^{\infty}_{n=1}e^{-m_{1}\alpha_{n}}$ is convergent for some $m_{1}\in\mathbb{N}$. It follows that there exists a $D>0$ such that
$\sum^{n}_{j=1}e^{p\alpha_{j}}\leq De^{(p+m_{1})\alpha_{n}}.$ (3.4)
Indeed, this is evident from the following estimates:
Then, we get
for every $n\in\mathbb{N}$.
@@
for every $k\in\mathbb{N}$. It follows that for every $k\in\mathbb{N}$, there exists a $D>0$ satisfying
$ne^{-\frac{1}{k}\alpha_{n}}\leq De^{-\frac{1}{2k}\alpha_{n}}$ (3.5)
for all $n\in\mathbb{N}$.
Since $\alpha$ is stable, there exists a $M\in\mathbb{N}$ such that $\alpha_{2t}\leq M\alpha_{t}$ for every $t\in\mathbb{N}$. Assume that $n=2t$ or $n=2t+1$ and $1\leq j\leq n$
$\alpha_{n}\leq M(\alpha_{n-j+1}+\alpha_{j})$ (3.6)
is satisfied for all $n\in\mathbb{N}$ and $j\leq n$.
@@
$\widecheck{T}_{\beta}\widecheck{T}_{\psi}=\widecheck{T}_{\psi}\widecheck{T}_{\beta}=\widecheck{T}_{\psi*\beta}.$
@@ 4. Topologizability and Power Boundedness of Convolution Operators on Power Series Spaces
Let $\theta\in\Lambda_{1}(\alpha)$. For every $p,n\in\mathbb{N}$, we have
that is,
@@
$\theta^{*k}:=\theta*\theta*\cdots*\theta$
@@
$\begin{split}\|\theta^{*k}\|_{p}&=\|\widehat{T}_{\theta^{*k}}e_{1}\|_{p}=\|\widehat{T}_{\theta}(\widehat{T}_{\theta^{*(k-1)}}e_{1})\|_{p}=\\ &\leq\|\theta\|_{2p}\|\widehat{T}_{\theta^{*(k-1)}}e_{1}\|_{2p}\leq\dots\leq\prod^{k}_{i=1}\|\theta\|_{2^{i}p}\end{split}$ (4.3)
@@
$\|\widehat{T}^{k}_{\theta}e_{n}\|_{p}\leq\left(\prod^{k}_{i=1}\|\theta\|_{2^{i}p}\right)\|e_{n}\|_{2p}.$ (4.4)
for every $p,k,n\in\mathbb{N}$. This ensures that $\widehat{T}_{\theta}$ is topologizable for every $\theta\in\Lambda_{1}(\alpha)$.
@@
The convolution operator $\widehat{T}_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ is topologizable for every $\theta\in\Lambda_{1}(\alpha)$. The operator $\widehat{T}_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ is m-topologizable if and only if for every $p\in\mathbb{N}$, there exists a $C_{p}>0$ satisfying
$\|\theta^{*k}\|_{p}\leq C^{k}_{p}$
for every $k\in\mathbb{N}$.
@@
Topologizability follows immediately from the inequality (4.2) and Lemma 2.5. If the operator $\widehat{T}_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ is m-topologizable, then for every $p\in\mathbb{N}$, there exist $q\in\mathbb{N}$ and $C_{p}>0$ such that
$\|\widehat{T}^{k}_{\theta}x\|_{p}\leq C^{k}_{p}\|x\|_{q}$
for every $k\in\mathbb{N}$ and $x\in\Lambda_{1}(\alpha)$. This gives us that for every $p\in\mathbb{N}$, there exists a $C_{p}>0$ such that
$\|\theta^{*k}\|_{p}=\|\widehat{T}^{k}_{\theta}e_{1}\|_{p}\leq C^{k}_{p}\|e_{1}\|_{q}\leq C^{k}_{p}$
for every $k\in\mathbb{N}$. The converse follows from the inequality (4.2). ∎
Given that the operators $\widehat{T}_{\theta}$ and $\widehat{T}_{\phi}$ commute and their product satisfies $\widehat{T}_{\theta}\widehat{T}_{\phi}=\widehat{T}_{\theta*\phi}$, we can conclude that their product is always topologizable for every $\theta,\phi\in\Lambda_{1}(\alpha)$.
@@
and if $j=2t+1\geq 2n$, we have $t+1\geq n$ and
Therefore, we have
$\alpha_{j}\leq M\alpha_{j-n+1}$
for all $j\geq 2n$. For every $p,n\in\mathbb{N}$, we have
that is,
@@
holds for every $p,n\in\mathbb{N}$. In Proposition 3.9, we showed that the product of two convolution operators, say $\widehat{T}_{\theta}$ and $\widehat{T}_{\phi}$, is again a convolution operator $\widehat{T}_{\theta*\phi}$. By using (4.5), we obtain
@@
$\begin{split}\|\theta^{*k}\|_{p}&=\|\widehat{T}_{\theta^{*k}}e_{1}\|_{p}=\|\widehat{T}_{\theta}(\widehat{T}_{\theta^{*(k-1)}}e_{1})\|_{p}\\ &\leq\|\theta\|_{Mp}\|\widehat{T}_{\theta^{*(k-1)}}e_{1}\|_{Mp}\leq\dots\leq\prod^{k}_{i=1}\|\theta\|_{M^{i}p}\end{split}$ (4.7)
@@
$\|\widehat{T}^{k}_{\theta}e_{n}\|_{p}\leq\left(\prod^{k}_{i=1}\|\theta\|_{M^{i}p}\right)\|e_{n}\|_{Mp}.$ (4.8)
holds for every $p,k,n\in\mathbb{N}$. This ensures that $\widehat{T}_{\theta}$ is topologizable for every $\theta\in\Lambda_{\infty}(\alpha)$.
@@
Let $\alpha$ be a stable sequence. The convolution operator $\widehat{T}_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\alpha)$ is topologizable for every $\theta\in\Lambda_{\infty}(\alpha)$. The operator $\widehat{T}_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\alpha)$ is m-topologizable if and only if for every $p\in\mathbb{N}$, there exists a $D_{p}>0$ satisfying
$\|\theta^{*k}\|_{p}\leq D^{k}_{p}$
for every $k\in\mathbb{N}$.
@@
Topologizability follows immediately from the inequality (4.6) and Lemma 2.5. If the operator $\widehat{T}_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\alpha)$ is m-topologizable, then for every $p\in\mathbb{N}$, there exist $q\in\mathbb{N}$ and $C_{p}>0$ such that
$\|\widehat{T}^{k}_{\theta}x\|_{p}\leq C^{k}_{p}\|x\|_{q}$
for every $k\in\mathbb{N}$ and $x\in\Lambda_{\infty}(\alpha)$. This gives us that for every $p\in\mathbb{N}$, we define $D_{p}:=C_{p}e^{q\alpha_{1}}$ such that
$\|\theta^{*k}\|_{p}=\|\widehat{T}^{k}_{\theta}e_{1}\|_{p}\leq C^{k}_{p}\|e_{1}\|_{q}=C^{k}_{p}e^{q\alpha_{1}}\leq D^{k}_{p}$
for every $k\in\mathbb{N}$. The converse follows from the inequality (4.6). ∎
We again note that the operators $\widehat{T}_{\theta}$ and $\widehat{T}_{\phi}$ commute and their product satisfies $\widehat{T}_{\theta}\widehat{T}_{\phi}=\widehat{T}_{\theta*\phi}$, then we can conclude that their product is always topologizable for every $\theta,\phi\in\Lambda_{\infty}(\alpha)$.